# SeTAR: Out-of-Distribution Detection with Selective Low-Rank Approximation

### v1

@@ 4.5 Analyses
SeTAR is effective on different model backbones.
To examine the effectiveness of SeTAR on different model backbones, we compare OOD detection performance with CLIP-large (CLIP ViT-L/14666https://huggingface.co/openai/clip-vit-large-patch14) and Swin Transformer (Swin-V2-B/16@256px-ImageNet1K [^7] https://huggingface.co/microsoft/swinv2-base-patch4-window16-256(Liu et al., 2021) in addition to CLIP-base. The Swin Transformer (Liu et al., 2022) is a novel unimodal model trained with the ImageNet1K classification dataset. We apply SeTAR to the image ViT as it lacks a text encoder. We follow common practice and consider two scoring functions for Swin Transformer: MSP (Hendrycks & Gimpel, 2017), using the softmax confidence score to detect OOD samples, and the Energy score (Liu et al., 2020), utilizing information in logits for OOD detection.
Table 8: Results for different model backbones.
@@
We set the temperature $T=0.1$ for the Energy score. As shown in Table 8, SeTAR outperforms baselines across all model backbones and scoring functions. Notably, applying SeTAR to Swin Transformer with the Energy score significantly reduces FPR95 by 20.61% on average. These results highlight SeTAR ’s potential for improving OOD detection with unimodal image encoders. Our results on SeTAR+FT (Table 2) further confirm its effectiveness across different model backbones.
@@ 5 Related Work
Previous work explores OOD detection with unimodal (DeVries & Taylor, 2018; Hendrycks & Gimpel, 2017; Hu & Khan, 2021; Zheng et al., 2020; Zhou et al., 2021) and multimodal (Fort et al., 2021; Ming et al., 2022; Tao et al., 2023; Miyai et al., 2023a) models. Numerous methodologies (Lee et al., 2018; Huang et al., 2021; Sun et al., 2022; Wang et al., 2022; Wu et al., 2023) have been developed to tackle OOD detection in computer vision. Existing CLIP-based OOD detection methods include zero-shot (Fort et al., 2021; Ming et al., 2022; Miyai et al., 2023b; Dai et al., 2023; Wang et al., 2023) and fine-tuning (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a). Zero-shot methods like MCM (Ming et al., 2022) and GL-MCM (Miyai et al., 2023b) don’t require in-distribution training data but may perform suboptimally due to domain gaps. Fine-tuning methods (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a) improve OOD detection by adapting to in-distribution data but risk damaging the pretraining representations, needing careful training strategies. Our proposed SeTAR complements various scoring functions and model backbones for OOD detection.
@@
Neural networks trained with over-parameterization often exhibit low-rank properties (Oymak et al., 2019). These properties are utilized in both model training (Povey et al., 2018; Hu et al., 2022) and post-hoc processing (Hajimolahoseini et al., 2021; Sharma et al., 2023). In training, some works (Sainath et al., 2013; Zhang et al., 2014; Zhao et al., 2016) impose low-rank constraints, while LoRA (Hu et al., 2022) adapts pretrained LLMs to downstream tasks using trainable low-rank matrices. For post-hoc processing, pruning methods (Yu et al., 2017; Hajimolahoseini et al., 2021) reduce weight matrix ranks by retaining top-K components from SVD. While pruning preserves model behavior, performance declines with increased intervention. Unlike these approaches, our work employs a selective rank reduction strategy.
@@ Appendix F More Detailed Experiment Results
In this section, we present additional detailed results from the main paper. This includes the detailed results of fine-tuned baselines on the ImageNet1K benchmark in Table 12; detailed ablation results on modality, $\mathrm{W_{p}}$, $\lambda$, and top-K in Table 13, Table 14, Table 17, and Table 19; and detailed results of SeTAR with different search algorithms, prune strategies and backbones in Table 16, Table 18 and Table 15.
@@
Table 15: Detail results for SeTAR with different backbones. denotes the result of our re-run.

### v2

@@ 4.5 Analyses
SeTAR is Effective on Different Architectures and Score Functions
We expand on Table 1 with results on ViT and CNN backbones and various score functions. For ViT-based models, we evaluate OOD detection using CLIP-large [^6] https://huggingface.co/openai/clip-vit-large-patch14 and Swin Transformer [^7] https://huggingface.co/microsoft/swinv2-base-patch4-window16-256 (Liu et al., 2021), alongside CLIP-base. The Swin Transformer (Liu et al., 2022) is trained on ImageNet1K. Since it lacks a text encoder, we apply SeTAR to the image ViT only. For Swin Transformer, we use two common scoring functions: MSP (Hendrycks & Gimpel, 2017), which leverages softmax confidence, and the Energy score (Liu et al., 2020), with $T=0.1$ for OOD detection. We also integrate CLIP-base
Table 8: Results for different ViT backbones.
@@
with the NegLabel score function (Jiang et al., 2024), which uses large-scale negative labels. As shown in Table 8, SeTAR consistently outperforms baselines across all backbones and scoring functions, significantly reducing FPR95 by relatively 20.61% with the Energy score on Swin Transformer. These results demonstrate SeTAR ’s effectiveness in improving OOD detection for unimodal image encoders, with further confirmation from SeTAR+FT results (Table 2) across different model backbones.
Table 9: Results on ResNet50. We use ImageNet1K as the ID dataset. is cited from Djurisic et al. (2023).
Method FPR↓ AUC↑ Method FPR↓ AUC↑
Softmax 66.95 81.99 ASH-P 50.32 89.04
Energy 58.41 86.17 ASH-B 22.73 95.06
ReAct 31.43 92.95 ASH-S 22.80 95.12
DICE 34.75 90.77 SeTAR 22.38 95.25
We further explore SeTAR’s potential on CNN architecture, and compare it with methods such as Softmax, Energy (Wu et al., 2023), ReAct (Sun et al., 2021), DICE (Sun & Li, 2022), and ASH (Djurisic et al., 2023) on ResNet50888https://download.pytorch.org/models/resnet50-19c8e357.pth. Since ResNet lacks local features for OOD loss, we conducted experiments using only ID loss. We applied low-rank approximation to the in- and out-feature dimensions of the convolutional layers, combined with ASH for search. As shown in Table 9, SeTAR establishes new state-of-the-art results on ResNet, demonstrating its effectiveness across both ViT and CNN architectures.
Near-OOD Results
Table 10: Near-OOD results on CLIP-base.
Method Category MCM Score GL-MCM Score
[…]
Vanilla Training-Free 89.28 63.88 85.62 67.63
SeTAR Training-Free 88.29 64.20 84.03 68.29
LoCoOp Training-Free 89.72 63.45 86.79 65.93
LoRA Finetuning 88.52 65.38 84.39 68.85
SeTAR+FT Finetuning 87.16 68.13 84.72 70.42
To further evaluate SeTAR’s performance on diverse OOD tasks, we test it on a more challenging near-OOD setting using ImageNet1K as the ID dataset and SSB-Hard (Vaze et al., 2022) as the OOD dataset. As shown in Table 10, SeTAR and SeTAR+FT outperform the baselines, demonstrating superior performance in near-OOD scenarios.
@@ 5 Related Work
Previous work explores OOD detection with unimodal (DeVries & Taylor, 2018; Hendrycks & Gimpel, 2017; Hu & Khan, 2021; Zheng et al., 2020; Zhou et al., 2021) and multimodal (Fort et al., 2021; Ming et al., 2022; Tao et al., 2023; Miyai et al., 2023a) models. Numerous methodologies (Lee et al., 2018; Huang et al., 2021; Sun et al., 2022; Wang et al., 2022; Wu et al., 2023) have been developed to tackle OOD detection in computer vision. Existing CLIP-based OOD detection methods include zero-shot (Fort et al., 2021; Ming et al., 2022; Miyai et al., 2023b; Dai et al., 2023; Wang et al., 2023; Jiang et al., 2024) and fine-tuning (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a). Zero-shot methods like MCM (Ming et al., 2022) and GL-MCM (Miyai et al., 2023b) don’t require in-distribution training data but may perform suboptimally due to domain gaps. Other approaches integrate external knowledge. For example, CLIPN (Wang et al., 2023) pre-trains a novel NO-encoder on the CC-3M dataset (Sharma et al., 2018) to empower CLIP’s "no" logic for zero-shot evaluation. NegLabel (Jiang et al., 2024) demonstrates better performance than CLIPN by introducing large-scale negative labels for enhanced label scoring. Fine-tuning methods (Ming & Li, 2023; Tao et al., 2023; Miyai et al., 2023a) improve OOD detection by adapting to in-distribution data but risk damaging the pretraining representations, needing careful training strategies. CNN-based OOD detection methods, including ReAct (Sun et al., 2021), ASH (Djurisic et al., 2023), DICE (Sun & Li, 2022), CIDER (Ming et al., 2023), PALM (Lu et al., 2024), and Hopfield Boosting (Hofmann et al., 2024), have also demonstrated strong results. However, methods like ReAct and ASH rely on the assumption that ID and OOD images produce distinct activations in models trained on ID data. This assumption does not hold in large-scale pre-trained models like CLIP, where activations for ID and OOD images are not significantly different, limiting the effectiveness of such approaches in enhancing CLIP’s zero-shot OOD detection capabilities. SeTAR, in contrast, offers high compatibility with various scoring functions (e.g., MCM, GL-MCM, MSP, Energy), multiple model backbones (e.g., CLIP, Swin, ResNet), and advanced OOD techniques such as NegLabel. Designed to be both lightweight and efficient, SeTAR addresses the demand for resource-efficient solutions in OOD detection.
@@
Neural networks trained with over-parameterization often exhibit low-rank properties (Oymak et al., 2019). These properties are utilized in both model training (Povey et al., 2018; Hu et al., 2022) and post-hoc processing (Hajimolahoseini et al., 2021; Sharma et al., 2023). In training, some works (Sainath et al., 2013; Zhang et al., 2014; Zhao et al., 2016) impose low-rank constraints, while LoRA (Hu et al., 2022) adapts pretrained LLMs to downstream tasks using trainable low-rank matrices. For post-hoc processing, pruning methods (Yu et al., 2017; Hajimolahoseini et al., 2021) reduce weight matrix ranks by retaining top-K components from SVD. While pruning preserves model behavior, performance declines with increased intervention. LASER (Sharma et al., 2023) focuses on pruning individual layers to enhance factual answering capabilities. It utilizes a simple greedy search strategy on a validation set, which is not applicable for OOD detection due to the absence of a validation set. In contrast, our approach introduces a selective rank reduction strategy specifically tailored for OOD detection. We systematically analyze and compare different greedy search techniques, evaluating their effectiveness across various layers and model backbones.
@@ 6 Conclusion
Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132, No. 62106138). We thank the anonymous reviewers for their insightful feedbacks on this work.
@@ Appendix F More Detailed Experiment Results
In this section, we present additional detailed results from the main paper. This includes the detailed results of fine-tuned baselines on the ImageNet1K benchmark in Table 14; detailed ablation results on modality, $\mathrm{W_{p}}$, $\lambda$, and top-K in Table 15, Table 16, Table 19, and Table 21; and detailed results of SeTAR with different search algorithms, prune strategies and backbones in Table 18, Table 20, Table 17 and Table 22.
@@
Table 17: Detail results for SeTAR with different backbones. is cited from Jiang et al. (2024). denotes the result of our re-run.
@@
Table 22: Detail results of ResNet50. We use ImageNet1K as the ID dataset. is cited from Djurisic et al. (2023).
Method iNaturalist SUN Places Texture Average
[…]
Softmax 54.99 87.74 70.83 80.86 73.99 79.76 68.00 79.61 66.95 81.99
Energy 55.72 89.95 59.26 85.89 64.92 82.86 53.72 85.99 58.41 86.17
ReAct 20.38 96.22 24.20 94.20 33.85 91.58 47.30 89.80 31.43 92.95
DICE 25.63 94.49 35.15 90.83 46.49 87.48 31.72 90.30 34.75 90.77
ASH-P 44.57 92.51 52.88 88.35 61.79 61.79 42.06 89.70 50.32 89.04
ASH-B 14.21 97.32 22.08 95.10 33.45 92.31 21.17 95.50 22.73 95.06
ASH-S 11.49 97.87 27.98 94.02 39.78 90.98 11.93 97.60 22.80 95.12
SeTAR 10.08 98.11 27.68 94.15 39.22 91.24 12.54 97.51 22.38 95.25